OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 16.

OpenStudy (calculusfunctions):

An equation for an ellipse is given as\[\frac{ x ^{2} }{ a ^{2} }+\frac{ y ^{2} }{ b ^{^{2}} }=1\]where 2a is the length of the major axis and 2b is the length of the minor axis.

OpenStudy (calculusfunctions):

Of course,the above equation presumes that the center of the ellipse as at the origin.

OpenStudy (anonymous):

so would x2/9 + x2/8 = 1 be the answer @calculusfunctions ??